OpenStudy (anonymous):

Simplify:4sqrt400/4sqrt5 Show your work.

OpenStudy (anonymous):

@Directrix @iGreen

OpenStudy (anonymous):

\[4\sqrt{400} /4\sqrt{5}\] Is this the question?

OpenStudy (anonymous):

yes

OpenStudy (mrnood):

400 = 4 * 100

OpenStudy (anonymous):

ok i understand that

OpenStudy (mrnood):

sqrt (a*b) = sqrt a * sqrt b so use the above to simplify sqrt (400)

OpenStudy (anonymous):

wait waht im counfuzzled how do you do that this is my first time doing a problem like this

OpenStudy (mrnood):

sqrt (400) = sqrt (4 * 100) sqrt (a*b) = sqrt a * sqrt b so use that to simplify sqrt (400)

OpenStudy (anonymous):

@iGreen what he mean lol

OpenStudy (mrnood):

There is nothing "lol" about this Are you going to do anything to wards this or are you waiting for the answer?

OpenStudy (anonymous):

yep if i understood it

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

@Lakeekee_

OpenStudy (anonymous):

help me please bae

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

make sure to show all the work you did please

OpenStudy (anonymous):

man

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

yo do you want help with this?

OpenStudy (anonymous):

@Tazmaniadevil do you want help with this?

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

do you want to learn how to do it or do you just want the answer?

OpenStudy (anonymous):

both lol

OpenStudy (anonymous):

lol respect, how bout learning how to do it, so you dont have to rely on ppl giving you the answer for the next 1000 problems you will have to do

OpenStudy (anonymous):

ok let geter done

OpenStudy (anonymous):

ok im going to draw it and then ask you questions step by step

OpenStudy (anonymous):

ok

OpenStudy (anonymous):